### v1

@@ Abstract
Survival prediction for cancer patients is critical for optimal treatment selection and patient management. Current patient survival prediction methods typically extract survival information from patients’ clinical record data or biological and imaging data. In practice, experienced clinicians can have a preliminary assessment of patients’ health status based on patients’ observable physical appearances, which are mainly facial features. However, such assessment is highly subjective. In this work, the efficacy of objectively capturing and using prognostic information contained in conventional portrait photographs using deep learning for survival predication purposes is investigated for the first time. A pre-trained StyleGAN2 model is fine-tuned on a custom dataset of our cancer patients’ photos to empower its generator with generative ability suitable for patients’ photos. The StyleGAN2 is then used to embed the photographs to its highly expressive latent space. Utilizing the state-of-the-art survival analysis models and based on StyleGAN’s latent space photo embeddings, this approach achieved a C-index of $0.677$, which is notably higher than chance and evidencing the prognostic value embedded in simple 2D facial images. In addition, thanks to StyleGAN’s interpretable latent space, our survival prediction model can be validated for relying on essential facial features, eliminating any biases from extraneous information like clothing or background. Moreover, a health attribute is obtained from regression coefficients, which has important potential value for patient care.
@@ 1 Introduction
Survival analysis for cancer patients is critical for optimal treatment selection and patient management. Patients with very adverse prognosis may need aggressive management and acute intervention or may not benefit from tumor-directed treatment at all. Patients with good prognostic constellation similarly may benefit from specific treatment approaches like dose escalation and radical treatment of all sites in radiotherapy. So far, a vast class of survival prediction models have been investigated, ranging from conventional regression algorithms [4, 13] to the latest radiomics [24, 3] and deep learning methods [27, 12, 14, 28]. With the successful applications in various fields including radiation oncology [6, 29, 7, 21], nowadays deep learning has become the main trend for patient survival prediction, because of its high representation power and ability to extract essential features from various types of data. Nevertheless, all the previous methods rely on data ranging from regular clinical records to commonly used medical imaging data (e.g., MRI and CT) as well as biological data (e.g., histologic and genetic information) [25]. Within radiotherapy practice, the acquisition of patient photographs is a routine measure implemented for identity verification purposes. Nevertheless, the potential utility in predicting patient survival and overall health status based on facial features derived from such widely available 2D portrait photos remains unexplored.
Facial features observable in 2D portrait photographs, like skin colorit and facial expression can reflect underlying pathophysiologic processes like muscle wasting, lowering of muscle tone or changes in microcirculation [19]. These features therefore hold considerable potential as indicators of a patient’s underlying health status. Many severe diseases will lead to evident facial changes, e.g., the most well-known diseases being genetic ones like the Down’s syndrome [5]. Assessing patients’ overall health and prognosis based on the facial expression has a long-standing history and dates back at least to hippocrates [30]. Clinicians with extensive experience frequently utilize patients’ physical appearance, as well as other related features, to estimate survival time and evaluate overall functional status. The Eastern Cooperative Oncology Group (ECOG) score is such a widely used metric for evaluating patient health status. This score is assigned by physicians and utilizes an ordinal 5-point scale to grade a patient’s overall health [19]. Hui et al. [8] were able to show that drooping of the nasolabial fold was a highly predictive indicator of imminent death within a three-day timeframe. Patient appearance has also been shown to be a factor in determining healthcare priority in emergency situations [2]. However, assessments made by human clinicians are inherently subjective and can therefore exhibit a high degree of intra- and inter-rater variance. In this work, we aim to achieve an objective assessment of cancer patients’ prognosis based on facial features using deep learning.
@@
Our contributions can be summarized as follows: a) To the best of our knowledge, this is the first attempt to use deep learning for cancer patient survival prediction based on 2D facial photographs; b) We propose a model that utilizes the StyleGAN expressive latent space feature representations for survival prediction; c) To provide interpretability and validation to the model, a “health” attribute is extracted to allow adjusting face photographs to either a healthier or more severe illness appearance. This manipulation has significant prognostic value for patient care and societal perception, underscoring its potential important clinical value.
@@
Figure 1: Exemplary face images and their corresponding synthetic images generated by StyleGAN without (w/o) and with fine-tuning (FT). The white rectangles are added for anonymization.
@@ 2.1 Data
The dataset utilized in this study comprises 16,657 2D portrait photographs of 13,523 individuals who have been diagnosed with cancer. Of this cohort, 52.5% are male and 47.4% are female. It is notable that 53.3% of the dataset is right censored, indicating that the time of death for these individuals is either unknown or has not yet occurred. In addition, the survival times for the non-censored patients included in the dataset ranged from 2 to 4,923 days. To mitigate potential biases that may be present in the photographs due to factors such as clothing and background, an automated face recognition process utilizing dlib [15] is employed to align and crop all images to a standardized size of $512\times 512$. The clinical information about the patients is retrieved, including international classification of diseases (ICD), gender, age, number of tumors, etc. Information acquired after the photo acquisitions is excluded.
@@ 2.2 Embedding Images to StyleGAN2 Latent Space
To accurately capture patients’ facial features from a given photo, a generative model should be able to accurately reproduce these expressive features into a generated image. A StyleGAN2 model pretrained on the FFHQ dataset [10] is fine-tuned using adaptive discriminator augmentation (ADA) [9] with our custom dataset of cancer patients’ face images. This step is necessary to increase StyleGAN’s generation ability for the special facial features in cancer patients and auxiliary medical equipment such as masks, breathing tubes and gauze. Fig. 1 shows the difference between synthesised images with and without fine-tuning. Without fine-tuning, the pretrained model fails to generate the mask and the bouffant cap in Fig. 1(b) and Fig. 1(e), respectively, while the fine-tuned model successfully re-synthesizes such structures.
@@ 2.3 Survival Analysis Models
To learn a more complex, non-linear relationship between the features for the hazard estimation, deep learning networks have been proposed, with DeepSurv [32, 12] being the state-of-the-art. The original DeepSurv method applies a multi-layer perceptron (MLP) to estimate the log-risk function $h(\boldsymbol{f})$ of the CoxPH model. To extract features from medical images, the MLP network can be replaced by a convolutional neural network (CNN) [32, 5]. In this work, a ResNet-18 and a VGG-16 network are investigated as examples of the CNN with image data, which are trained in an end-to-end approach. As CNN-extracted features are highly abstract, in this work, we propose to integrate the StyleGAN network to the DeepSurv framwork for the prediction of continuous survival time, i.e., utilizing the expressive StyleGAN photo embedding $\boldsymbol{z}$ as the input features of the CoxPH and DeepSurv models. Our method uses a two-step training with the benefit of good explainability. The first step provides a comprehensive set of features, thanks to the well-established StyleGAN latent space. The second step selects the features from this set, which are very important for survival prognosis. Such features can be later visualized directly by the existing StyleGAN generator.
@@ 2.4 Latent Space Manipulation
In this work we propose two tests to visually and qualitatively evaluate the semantic meaning of our photo embedding as well as the prognostic value it holds for the task of survival prediction. Each of the tests can be conducted by simple manipulations of the latent vector $\boldsymbol{z}$.
Synergetic Effect of Joint Features: Leveraging the linearity of the CoxPH regression model, we fitted the $512$-dimensional feature vectors $\boldsymbol{z}$ to the known survival data,
@@
2.5 Experimental Parameters
StyleGAN Fine-Tuning: The StyleGAN version $2$ was utilized and fine-tuned using transfer learning from the FFHQ dataset of human faces, which had a resolution of $512\times 512$. The fine-tuning process took $88$ hours and resulted in a final FID score of $5.1$ after $3200$k images. The fine-tuning was performed on a single NVIDIA Quadro RTX $8000$ $48$GB GPU.
Iterative Embedding Optimization: The optimization process was initialized with the latent vector $\overline{\boldsymbol{z}}$ of the average face and performed $800$ gradient descent steps for the optimization using Adam optimizer with an initial learning rate of $0.01$, $\beta_{1}=0.9,\beta_{2}=0.999,\text{ and }\epsilon=1e^{-8}$
DeepSurv: In this work, the DeepSurv model was trained with a train-test split of 80% for training and validation and 20% for testing using a batch size of $128$ and with the Adam optimizer and a learning rate of $0.001$. The network architecture consisted of a MLP with $6$ layers with batch normalization and dropout of $0.4$ to enhance the model’s generalization and prevent overfitting. Furthermore, the Kaplan-Meier method was employed to estimate the censoring distribution, which is crucial for survival analysis tasks.
Data Anonymization Disclaimer: The data presented in this work has been anonymized to protect the privacy of individuals. Any personal information has been removed or de-identified, and the data cannot be used to identify any specific individuals. Facial images are either synthetic, anonymized, or projections that have been reviewed by clinical experts to ensure that no identification of individuals is possible.
Table 1: Survival prediction results of different methods using DeepSurv (DS) and CoxPH models. C-index and Brier score (B-Score) is caculated for all models.
@@ 2.5 Experimental Parameters
Embeddings DS+MLP 0.677 0.130 Clinic. Data CoxPH 0.690 0.151
@@
3 Results and Discussion
3.1 Survival Prediction Results
The survival prediction results are displayed in Table 1. Our approach for continuous survival time prediction using DeepSurv based on StyleGAN photo embeddings achieved a mean concordance index (C-index) of $0.677$ and a mean Brier Score of $0.130$. To put this into perspective, the model’s survival predictions are evidently higher than chance. This reflects that the model uses key survival features from relevant prognostic facial attributes embedded in the latent vectors. As also shown in Table 1, the DeepSurv model using patient’s portraits fed directly through a CNN achieved a C-index of $0.47$ and $0.51$ using the ResNet-18 and the VGG-16 networks respectively. In this case, the CNNs failed to extract the needed prognostic information directly from the portrait photos.
As also shown in Table 1, survival prediction models based on conventional clinical features achieved a C-index of 0.691 and 0.729 using CoxPH and DeepSurv models respectively, which is comparable to what was achieved by the embeddings of simple 2D portraits. Combining clinical features with photo embeddings boosted the C-index to 0.787 highlighting again the complementary prognostic value of StyleGAN’s embeddings. It is worth mentioning that the combination was done by early fusion, in which the inputs were simply concatenated before feeding them into the network. A more advanced fusion method could potentially produce better results, which is our future work.
@@ 3.1 Survival Prediction Results
$\beta=5$
$\beta=10$
$\beta=20$
Figure 3: Manipulating the health attribute with different $\beta$ values for regenerated patients’ images. The synthesized image without manipulation is at $\beta=0$ (third column). These are projections that do not allow identification of real persons
3.2 Model Interpretability with “Health” Attribute
Fig. 3 demonstrates the synergetic effect of manipulating patient’s images in the latent space using the health attribute (Eqn. (4)). In Fig. 3 the appearances of the patients are controlled by $\beta$. With a negative value for $\beta$, the patients appear to be healthier with a reddish skin colorit, full cheeks and a smiling expression with a strong tone of the facial muscles. In contrast, with a positive value for $\beta$, there is loss of masticatory muscles and subcutaneous fat with a pale skin colorit. Moreover, the tone of the facial muscles is reduced with drooping of the nasolabial fold, an overall more relaxed appearance of facial features and an increased “tippiness” of the nose. These features are consistent with long-standing clinical descriptions of prognostically relevant changes in facial appearance [8, 30]. Most importantly, the change is mainly located in the head region instead of their clothes or the background. The model is also able to reflect patients’ mental status to some degree by extracting their facial expressions.
@@ 3.2 Model Interpretability with “Health” Attribute
An observation that could not be neglected however is the similarity between progressing in age and worse prognostic features, as age is a common risk indicator. While one can assume an underlying correlation between age and general health related facial features, we addressed a potential bias of the model to only rely on age features for the prognosis prediction task. At the top row of Fig. 3, it is clearly observable that when editing the patient’s image based on the health attribute, distinct facial features are affected compared to when editing based on the age attribute, as shown at the top row of Fig. 4.
Based on each feature’s importance for the task of survival prediction, a evident change is observed in the manipulated image. The images in the second row of Fig. 4 that changing random features in the latent space would barely change features with any prognostic value in a meaningful way. This demonstrates that, as in Fig. 3, the StyleGAN regression model is indeed using meaningful facial features instead of bias or random information for survival prediction.
@@
$\beta=5$
$\beta=10$
$\beta=20$
Figure 4: Exemplary patients’ images generated by manipulating the age attribute (top row), and a single feature at random dimension $132$ (bottom row). These are projections that do not allow identification of real persons.
Using a dataset with only cancer patients was intentional, as the model is specifically designed for survival prediction of cancer patients, not for the general population. Nevertheless, the survival time in our dataset ranges from 3 days to 13 years, covering a diverse range of health statuses. Compared with the short-survival patients, long-survival patients are relatively healthier. Many of them have a long survival time just close to people without cancer. Therefore, health features can still be extracted.
In this preliminary work, we intentionally decided not to split the dataset into multiple subgroups, as the clinical literature indicated that facial prognostic features are universal and transcend cancer types. Moreover, splitting solely by tumor type also would not have addressed additional clinical dimensions like disease stage, specific molecular subtype etc. and was expected to greatly reduce the generalizability of the model. Therefore, we included images that were captured when the patient was admitted for radiotherapy at various disease stages.
Currently, the well-established cancer prognosis standard is the AJCC TNM staging [1], which incorporates tumor, lymph nodes, and metastasis as the primary cues. Our model provides an additional tool for prognosis estimation, which is less influenced by human’s wishful thinking, complementing traditional clinical variables. Consequently, we showed a significant improvement in prognostic accuracy over relying solely on clinical variables, thereby supporting more effective oncologic decision-making.

### v3

@@ Abstract
Survival prediction for cancer patients is critical for optimal treatment selection and patient management. Current patient survival prediction methods typically extract survival information from patients’ clinical record data or biological and imaging data. In practice, experienced clinicians can have a preliminary assessment of patients’ health status based on patients’ observable physical appearances, which are mainly facial features. However, such assessment is highly subjective. In this work, the efficacy of objectively capturing and using prognostic information contained in conventional portrait photographs using deep learning for survival prediction purposes is investigated for the first time. A pre-trained StyleGAN2 model is fine-tuned on a custom dataset of our cancer patients’ photos to empower its generator with generative ability suitable for patients’ photos. The StyleGAN2 is then used to embed the photographs to its highly expressive latent space. Utilizing state-of-the-art survival analysis models and StyleGAN’s latent space embeddings, this approach predicts the overall survival for single as well as pan-cancer, achieving a C-index of $0.680$ in a pan-cancer analysis, showcasing the prognostic value embedded in simple 2D facial images. In addition, thanks to StyleGAN’s interpretable latent space, our survival prediction model can be validated for relying on essential facial features, eliminating any biases from extraneous information like clothing or background. Moreover, our approach provides a novel health attribute obtained from StyleGAN’s extracted features, allowing the modification of face photographs to either a healthier or more severe illness appearance, which has significant prognostic value for patient care and societal perception, underscoring its potential important clinical value.
@@ 1 Introduction
Survival analysis for cancer patients is critical for optimal treatment selection and patient management. Patients with very adverse prognosis may need aggressive management and acute intervention. However, patients with good prognostic constellations similarly may benefit from specific treatment approaches like dose escalation and radical treatment of all sites in radiotherapy. So far, a vast class of survival prediction models have been investigated, ranging from conventional regression algorithms [5, 12] to the latest radiomics [22, 4] and deep learning methods [25, 11, 13, 26]. With the successful applications in various fields, nowadays deep learning has become the main trend for patient survival prediction, because of its high representation power and ability to extract essential features from various types of data. Nevertheless, all the previous methods rely on data ranging from regular clinical records to commonly used medical imaging data (e.g., MRI and CT) as well as biological data (e.g., histologic and genetic information) [23]. Within radiotherapy practice, the acquisition of patient photographs is a routine measure implemented for identity verification purposes. Nevertheless, the potential utility in predicting patient survival and overall health status based on facial features derived from such widely available 2D portrait photos remains unexplored.
Facial features observable in 2D portrait photographs, like skin colorit and facial expression can reflect underlying pathophysiologic processes like muscle wasting, lowering of muscle tone or changes in microcirculation [18]. These features therefore hold considerable potential as indicators of a patient’s underlying health status. Many severe diseases will lead to evident facial changes, e.g., the most well-known diseases being genetic ones like the Down’s syndrome [6]. Clinicians with extensive experience frequently utilize patients’ physical appearance, as well as other related features, to estimate survival time and evaluate overall functional status. The Eastern Cooperative Oncology Group (ECOG) score is such a widely used metric for evaluating patient health status [18]. Hui et al. [7] were able to show that drooping of the nasolabial fold was a highly predictive indicator of imminent death within a three-day timeframe. Patient appearance has also been shown to be a factor in determining healthcare priority in emergency situations [1]. However, assessments made by clinicians are inherently subjective and can therefore exhibit a high degree of intra- and inter-rater variance. In this work, we aim to achieve an objective assessment of cancer patients’ prognosis based on facial features using deep learning.
@@
Our contributions can be summarized as follows: a) To the best of our knowledge, this is the first attempt to use deep learning for cancer patient survival prediction based on 2D facial photographs; b) We propose a model that utilizes the StyleGAN expressive latent space feature representations for survival prediction. c) We introduce a novel addition to traditional prognostic methods by using StyleGAN embeddings alongside clinical parameters to improve survival prediction accuracy. d) To provide interpretability and validation to the model, a “health" attribute is extracted to allow adjusting face photographs to either a healthier or more severe illness appearance. This manipulation has significant prognostic value for patient care and societal perception, underscoring its potential important clinical value.
@@
Figure 1: Exemplary anonymized face images and their corresponding synthetic images generated by StyleGAN without (w/o) and with fine-tuning (FT).
@@ 2.1 Data
The dataset utilized in this work comprises 13,503 2D portrait photographs captured during the admission of individuals who have been diagnosed with cancer. Of this cohort, 52.5% are male and 47.5% are female. Notably, 57.3% of the dataset is right censored. In addition, the survival times for the non-censored patients range from 2 to 4,791 days. To mitigate potential biases in the photographs due to factors such as clothing and background, an automated face recognition process utilizing dlib [14] is employed to align and crop all images to a standardized size of $512\times 512$. The clinical information about the patients is retrieved, including international classification of diseases (ICD), gender, age, number of tumors, etc. Information acquired after the photo acquisitions is excluded.
Data Anonymization Disclaimer: This work’s data is anonymized to safeguard individual privacy. Facial images are synthetic, anonymized, or expert-reviewed projections, ensuring individuals cannot be identified.
@@ 2.2 Embedding Images to StyleGAN2 Latent Space
To accurately capture patients’ facial features from a given photo, a generative model should be able to accurately reproduce these expressive features into a generated image. A StyleGAN2 model pretrained on the FFHQ dataset [9] is fine-tuned using adaptive discriminator augmentation (ADA) [8] with our custom dataset of cancer patients’ face images. This step is necessary to increase StyleGAN’s generation ability for the special facial features in cancer patients and auxiliary medical equipment such as masks, breathing tubes and gauze. Fig. 1 shows the difference between synthesised images with and without fine-tuning. Without fine-tuning, the pretrained model fails to generate the mask and the bouffant cap in Fig. 1(b) and Fig. 1(e), respectively.
@@ 2.3 Survival Analysis Models
To learn a more complex, non-linear relationship between the features for the hazard estimation, deep learning networks have been proposed, with DeepSurv [29, 11] being the state-of-the-art. The DeepSurv method applies a multi-layer perceptron (MLP) to estimate the log-risk function $h(\boldsymbol{f})$ of the CoxPH model. To extract features from medical images, the MLP network can be replaced by a convolutional neural network (CNN) [29, 6]. In this work, a ResNet-18 and a VGG-16 network are investigated as examples of the CNN with image data. As CNN-extracted features are highly abstract, in this work, we propose to integrate the StyleGAN network to the DeepSurv framework for the prediction of continuous survival time, i.e., utilizing the expressive StyleGAN photo embedding $\boldsymbol{z}$ as the input features of the CoxPH and DeepSurv models.
@@ 2.4 Latent Space Manipulation
In this work, we propose two tests to visually and qualitatively evaluate the semantic meaning of our photo embedding as well as the prognostic value it holds for the task of survival prediction. Each of the tests can be conducted by simple manipulations of the latent vector $\boldsymbol{z}$.
Health Attribute Manipulation: Leveraging the linearity of the CoxPH regression model, we fit the $512$-dimensional vectors $\boldsymbol{z}$ to the known survival data,
@@
3 Results and Discussion
3.1 Pan-Cancer Analysis for Survival Prediction Results
The survival prediction results are displayed in Table 1. Our approach for continuous survival time prediction using DeepSurv based on StyleGAN photo embeddings achieved for pan-cancer analysis a mean concordance index (C-index) of $0.680$ and a mean Brier Score of $0.200$. To put this into perspective, the model’s survival predictions are evidently higher than chance. This reflects that the model uses key survival features from relevant prognostic facial attributes embedded in the latent vectors. As also shown in Table 1, the DeepSurv model using patients’ portraits fed directly through a CNN achieved a C-index of $0.51$ and $0.47$ using the ResNet-18 and the VGG-16 networks respectively. In this case, the CNNs failed to extract the needed prognostic information directly from the portrait photos. It is worth noting that the comparison methods (i.e., DS+ResNet and DS+VGG with input photo data) used end-to-end training instead of two-step training as our proposed method (i.e., DS+MLP with input combined embedding and clinical data).
As also shown in Table 1, survival prediction models based on conventional clinical features achieved a C-index of 0.690 and 0.729 using CoxPH and DeepSurv models respectively, which is comparable to what was achieved by the embeddings of simple 2D portraits. Combining clinical features with photo embeddings using late fusion boosted the C-index to 0.787 highlighting again the complementary prognostic value of StyleGAN’s embeddings.
Table 1: Survival prediction results of different methods using DeepSurv (DS) and CoxPH models. C-index and Brier score (B-Score) are calculated for all models.
@@ 3.1 Pan-Cancer Analysis for Survival Prediction Results
Embeddings DS+MLP 0.680 0.200 Clinic. Data CoxPH 0.690 0.151
@@
Figure 3: A subgroup analysis on 16 cancer types with corresponding single C-indices. The dashed line represents the average C-index of the pan-cancer analysis.
3.2 Subgroup Analysis for Survival Prediction Results
Intriguingly, as shown in Fig  3, the developed pan-cancer model showed good prognostic accuracy for nearly all cancer sub-types, suggesting that prognostically-relevant facial features are present in most malignancies. However, there were also notable differences in observed prognostic accuracy. High prognostic accuracy was observed for endocrine malignancies (including small-cell cancers), metastatic cancer of unknown primary (CUP), urothelial cancer, breast cancer and melanoma, with C-indices of 0.786, 0.738, 0.723, 0.697 and 0.678, respectively, whereas notably lower accuracy was observed for lower and upper gastrointestinal cancers, prostate cancer, hematologic malignancies, sarcoma and especially renal cancer, with C-indices of 0.609, 0.595, 0.639, 0.638, 0.631 and 0.52, respectively. From a clinical perspective, this pattern could be derived from differences in tumor biology, as hematologic malignancies as well as sarcomas have a different origin and a distinct biology from carcinomas [2]. In addition, prostate cancer, colorectal cancer, and renal cancer are slowly proliferating tumors with markedly slower disease progression, lower tendency for metastasis and good systemic treatment options that typically result in patients showing less severe systemic signs of the disease [3].
@@ 3.2 Subgroup Analysis for Survival Prediction Results
Figure 4: Manipulating the health attribute with different $\beta$ values for regenerated patients’ images. The synthesized image without manipulation is at $\beta=0$ (third column).
3.3 Model Interpretability with Health Attribute
Fig. 4 demonstrates the synergetic effect of manipulating patient’s images in the latent space using the health attribute (Eqn. (4)). In Fig. 4 the appearances of the patients are controlled by $\beta$. With a negative value for $\beta$, the patients appear to be healthier with a reddish skin colorit, full cheeks, a smiling expression and a strong tone of the facial muscles. In contrast, with a positive value for $\beta$, there is a loss of masticatory muscles and subcutaneous fat with a pale skin colorit. Moreover, the tone of the facial muscles is reduced with drooping of the nasolabial fold, an overall more relaxed appearance of facial features and an increased “tippiness" of the nose. These features are consistent with long-standing clinical descriptions of prognostically relevant changes in facial appearance [7, 27]. Most importantly, the change is mainly located in the head region instead of their clothes or the background. The model is also able to reflect patients’ mental status to some degree by extracting their facial expressions.
@@ 3.3 Model Interpretability with Health Attribute
An observation that could not be neglected however is the similarity between progressing in age and worse prognostic features, as age is a common risk indicator. While one can assume an underlying correlation between age and general health related facial features, we addressed a potential bias of the model to only rely on age features for the prognosis prediction task. As shown in Fig. 4, it is clearly observable that when editing the patient’s image based on the health attribute, distinct facial features are affected compared to when editing based on the age attribute, as shown in Fig. 5.
Based on the feature importance indicated by the regression model’s coefficients $\boldsymbol{w}$ in Eqn.(4), changes in the embedding vector elements with corresponding higher coefficients have a more observable impact on the attributes of the generated images. This demonstrates that, as in Fig. 4, the StyleGAN regression model is indeed using meaningful facial features instead of bias or random information for survival prediction.
@@
Figure 5: Exemplary patients’ images generated by manipulating the age attribute.